User Tools

Site Tools


handover:a1517

Achtung! R1 TRANSFER IN PROGRESS! MONITORING REQUIRED!

R1 monitoring instructions: operations:monitoring.ivs.etransfer

13:18 UT: transfer seems to be finished (Lucia)

A1517

system checks done fro all three antennas at the end of the experiment. Records included in the a1517 logs. Monica data exported to a1517$s_check.txt/png. (Lucia)

Hobart 12m:

Disk VSN: HOB+0122/32000

Data volume at beginning: 22157.891 GB

  • Missed scans 154-1730 through 154-1732a due to late finish of previous experiment (Ross).

UT 01:07 Disk deficit of 50GB (JS)

UT 02:38 Disk deficit of 68GB (JS)

UT 05:31 Disk deficit of 93GB (JS)

UT 13:05 disk deficit 150 GB (Lucia)

Katherine 12m:

Disk VSN:HOB+0125/32000

Data volume at beginning: 4201.143 GB

  • Missed scans 154-1730 through 154-1732b due to late finish of previous experiment (Ross).
  • 16:30 UT: Maser shows blue number! AHM power ~/+ dropped to 23.7 - Hobart has 24.2 and is green. Keep watching! (Lucia)

Yarragadee 12m:

Disk VSN:HOB+0129/32000

Data volume at beginning: 6270.240 GB

  • Missed scans 154-1730 through 154-1732b due to late finish of previous experiment (Ross).
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/a1517.txt · Last modified: 2015/06/04 17:54 by Jim Lovell